A football team has won 54 out 75 games played. With 45 games still to go, what percentage of these must be won to give the team a record for the season of 60% games won out of all games played? ( answer is 40% show your solution )
Answer by richwmiller(17219)

(54+x)/(75+45)=(54+x)/(120)=.60
x=18
18/45=2/5=.40=40%